OpenStudy (anonymous):

use logarithms to solve the exponential equation show your work 1 = 3.2 X 10^x + 0.68

OpenStudy (anonymous):

"X" means "multiply"?

OpenStudy (anonymous):

1) 0.32 = 3.2 * 10 ^ x 2) 0.1 = 10 ^ x 3) 10 ^ (-1) = 10 ^ x so x = -1, and u don't need 2 use logs